OpenStudy (vayacondios1):

in triangle PQR, angle P and angle Q are complimentary angles. if sin Q=4/5, cos P+cos Q=? no idea how to work it will fan all that help

jhonyy9 (jhonyy9):

do you know what mean complementary angles ?

OpenStudy (vayacondios1):

yes

jhonyy9 (jhonyy9):

so than what mean complementary angles ?

OpenStudy (vayacondios1):

they add up to 90

jhonyy9 (jhonyy9):

yes right so than you drawn this triangle how will be ?

OpenStudy (vayacondios1):

3 30 degree angles?

OpenStudy (vayacondios1):

acute

jhonyy9 (jhonyy9):

so you know that angle R has a measure of 90 degree - yes ?

OpenStudy (vayacondios1):

yes

jhonyy9 (jhonyy9):

bc. angle Q and angle P make together 90 degree

OpenStudy (vayacondios1):

oh so it equals 90

jhonyy9 (jhonyy9):



jhonyy9 (jhonyy9):

hope this will help you

OpenStudy (vayacondios1):

ok

jhonyy9 (jhonyy9):

using pythagora"s you can calculi easy the length of RQ - yes ?

OpenStudy (vayacondios1):

a2+b2=c2 right

jhonyy9 (jhonyy9):

yes

OpenStudy (vayacondios1):

i got 6.4 for the missing side

jhonyy9 (jhonyy9):

not is correct 16 + x^2 = 25 x^2 = 25-16 x^2 = 9 x= 3

jhonyy9 (jhonyy9):

yes ?

OpenStudy (vayacondios1):

oh i should have know since it was bigger than the hypotenuse

jhonyy9 (jhonyy9):

yes so now you need to calculi these cos of angle P and angle Q - ok. ?

OpenStudy (vayacondios1):

ok thank you

jhonyy9 (jhonyy9):

do you know what say the cos law ?

OpenStudy (vayacondios1):

cos q i got 3/5 and cos p i got 4/5

OpenStudy (vayacondios1):

so 1.4

OpenStudy (vayacondios1):

ow nm thats wrong

jhonyy9 (jhonyy9):

cos q and cos p are right sure

jhonyy9 (jhonyy9):

so than now cos p + cos q = ?

OpenStudy (vayacondios1):

how do you type it in a calculator

jhonyy9 (jhonyy9):

3/5 +4/5 = ?

OpenStudy (vayacondios1):

1.4

jhonyy9 (jhonyy9):

dont need to use calculator

OpenStudy (vayacondios1):

ok

jhonyy9 (jhonyy9):

3/5 +4/5= 7/5 = 1,4

jhonyy9 (jhonyy9):

ok.?

OpenStudy (vayacondios1):

how many degrees is the angle?

jhonyy9 (jhonyy9):

what angle ?

OpenStudy (vayacondios1):

oh nm sorry

OpenStudy (vayacondios1):

thank you so much you were big help

jhonyy9 (jhonyy9):

was my pleasure anytime good luck bye bye

OpenStudy (vayacondios1):

bye